# Toward Vu’s conjecture

### v1

@@ 1.2. Vu’s conjecture
In a graph $G$, the codegree of two vertices $u,v$ is the number of distinct neighbors that are common to both $u$ and $v$. Given $\zeta>0$, how large can the chromatic number be in a graph with maximum degree at most $\Delta$ and maximum codegree at most $\zeta\Delta$? Because of its connection to Kahn’s important result on the list chromatic index of linear hypergraphs [39] and in turn to the well-known Erdős–Faber–Lovász conjecture [40, 31, 41], Vu [64] famously proposed the following bold conjecture in 2002.
@@
This conjecture remains wide open even if $\chi_{\ell}(G)$ is replaced by the ordinary chromatic number $\chi(G)$. Moreover, even the much weaker statement that $G$ satisfies $\alpha(G)\geq(1/\zeta-\varepsilon)(n/\Delta)$ is still open and was also conjectured by Vu [64]. As mentioned earlier, Conjecture 1.1 is a far-reaching generalization of a theorem of Kahn [39] on the list chromatic index of linear hypergraphs. Vu [64, p. 109] emphasized its significance, noting that: “The bound in [Conjecture 1.1], if true, would be an amazing result. For instance, it would immediately imply a deep theorem of Kahn on the list chromatic index of hypergraphs.” Kahn’s theorem, in turn, implies the asymptotic form of the Erdős–Faber–Lovász conjecture [40, 31, 41]. Recently, Kelly, Kühn, and Osthus [43] confirmed a special case of Conjecture 1.1 that also recovers several of its significant consequences.
@@ 1.3. Proof overview
Key to our analysis is a new concentration tool — a variant of Talagrand’s inequality designed to handle exceptional outcomes — combined with a Turán-type argument to control vertices of high degree; see also Lemma 2.3 and the discussion afterwards. Concentration tools of this type have been used extensively in recent years [47, 22, 10, 12, 11], so we believe that our approach may be of independent interest.
Our proof of the concentration of $|X^{\prime}|$ proceeds as follows. Rather than attempting to establish concentration of $|X^{\prime}|$ directly, we first obtain concentration for two auxiliary random variables $|X^{\prime}_{1}|$ and $|X^{\prime}_{2}|$, where $X^{\prime}_{1}$ is the number of vertices $w\in N_{L}(u,c)\cap N_{L}(v,c)$ satisfying a, and $X^{\prime}_{2}$ is the number of vertices $w\in N_{L}(u,c)\cap N_{L}(v,c)$ satisfying a but not b. Since $|X^{\prime}|=|X^{\prime}_{1}|-|X^{\prime}_{2}|$, this reduction allows us to deduce concentration of $|X^{\prime}|$. The following conditions are crucial for bounding the parameter $\beta$ in our version of Talagrand’s inequality (Lemma 2.3); $\beta$ determines a Lipschitz-like property for a random variable.
@@
    Let $\Omega^{*}$ be the exceptional event that there exists a color $c^{\prime}$ such that “many” vertices in $X$ are activated and assigned the color $c^{\prime}$. Our concentration inequality allows us to assume that this exceptional event does not occur.
@@ 1.4. Algorithmic considerations
The rest of the paper is organized as follows. In Section 2, we collect some preliminary facts and the main tools we need. In Section 3, we derive Theorem 1.2 and Corollary 1.3 from Theorem 1.4. In Section 4, we describe the Wasteful Coloring Procedure formally and introduce the key “nibble lemma” at the heart of the procedure. In Section 5, we prove our version of Talagrand’s inequality to handle exceptional outcomes, which we employ in the proof of the nibble lemma in Section 6. Finally, in Section 7, we iteratively apply the nibble lemma to prove Theorem 1.5.
@@ 2. Preliminaries
In this section, we outline prior known results which are necessary for our proofs.
@@
Suppose that $G$ is a bipartite graph with parts $A$ and $B$ of sizes $m$ and $n$, respectively. Suppose that $G$ has no subset $X\subseteq A$ of size $t$ and no subset $Y\subseteq B$ of size $s$ for which $G[X\cup Y]$ is a complete bipartite graph. Then,
@@
The following is a straightforward corollary of Lemma 2.1, and we provide a proof for completeness.
@@
Let $G$ be a graph with vertex subsets $A$ and $B$ of sizes $m$ and $n$, respectively, for which $A\cup B=V(G)$ and every edge of $G$ has an endpoint in $A$ and an endpoint in $B$. Suppose that $G$ has no subset $X\subseteq A$ of size $t$ and $Y\subseteq B$ of size $s$ for which $X\cap Y=\emptyset$ and $[X,Y]$ is a $K_{s,t}$. Then,
@@
Let $H$ be a bipartite graph constructed as follows. Let $H$ have disjoint partite sets $A^{\prime}$ and $B^{\prime}$. For each vertex $a\in A$, add a vertex $a^{\prime}$ to $A^{\prime}$, and for each $b\in B$, add a vertex $b^{\prime}$ to $B^{\prime}$. For each pair $(a,b)\in A\times B$, if $ab\in E(G)$, then add an edge $a^{\prime}b^{\prime}$ to $E()$. Since each edge of $G$ corresponds to at least one edge of $H$, clearly $e(H)\geq e(G)$. Furthermore, clearly $H$ is $K_{s,t}$-free. Therefore, by Lemma 2.1, we have
$e(G)\leq e(H)<(t-1)^{1/s}nm^{1-1/s}+(s-1).$
@@
Next, as mentioned in Section 1.3, we develop a concentration inequality that is similar to that of Mahdian [49] but can handle exceptional outcomes; see Section 5 for its proof.
Lemma 2.3.
Let $X_{1},\dots,X_{m}$ be a set of independent random variables (often called trials), and let $(\Omega,\Sigma,\Pr)$ be the product space describing all random outcomes of $X_{1},\dots,X_{m}$. Let $\Omega^{*}\subseteq\Omega$ be some set of exceptional outcomes. Let $R_{1},\ldots,R_{n}$ be indicator random variables, each defined as a function of the trials $X_{1},\ldots,X_{m}$, and let $R=R_{1}+\dots+R_{n}$. Suppose that for each non-exceptional outcome $\mathbf{x}=(x_{1},\dots,x_{m})\in\Omega\setminus\Omega^{*}$ of the trials, the following hold for some $\beta,D>0$:
@@
  • $\sum_{i=1}^{n}|W_{i}(\mathbf{x})|\leq D$.
  • $\Pr(\Omega^{*})\leq 1/6$.
Then, for all $\tau\geq 12\sqrt{\pi\beta D}+2\Pr(\Omega^{*})\sup_{\mathbf{x}\in\Omega^{*}}R(\mathbf{x})$,
$\Pr\left(|R-\mathbb{E}[R]|\geq\tau\right)<6\exp\left(-\frac{\tau^{2}}{24\beta D}\right)+\Pr(\Omega^{*}).$
We note that by letting $\Omega^{*}=\emptyset$ above, we obtain a similar concentration inequality to that of Mahdian [49]. Informally, we apply Lemma 2.3 as follows. Recall the informal description of the Wasteful Coloring Procedure on page 1.3. For each vertex $v$, let $A_{v}$ be the random variable such that $A_{v}=0$ if $v$ is not activated and $A_{v}=\phi(v)\in L(v)$ otherwise. We let $\Omega$ be the product space over the independent random variables $\{A_{v}\}_{v\in V}$. The set $\Omega^{*}\subseteq\Omega$ of exceptional events that we use depends on the random variable that we are interested in concentrating. Typically, we let $\Omega^{*}$ consist of outcomes in which, for a fixed vertex $v$, there is some color $c$ such that many vertices $u\in N(v)$ satisfy $A_{u}=c$. This allows us to control the value of the parameter $\beta$ in Lemma 2.3. The exact application of Lemma 2.3 differs very slightly from the above due to the introduction of equalizing coin flips (see Section 4).
@@
For every $s\geq 2$, there exists a constant $C$ such that the following holds for all $\zeta\in[\log^{-16s}\Delta,1]$. If $G$ is a graph with maximum degree $\Delta\geq 1$ and $s$-codegree at most $\zeta\Delta$, then $\chi(G)\leq C\zeta^{1/16s}\Delta$.
@@ 4. The Wasteful Coloring Procedure
5. An exceptional variant of a concentration inequality
In this section, we prove Lemma 2.3-an exceptional version of the concentration inequality used by Mahdian in [49] to prove an upper bound on the strong chromatic index of $C_{4}$-free graphs. While the concentration inequality of Li and Postle [47, Theorem 2.4] (originally appearing in a manuscript of Delcourt and Postle [27]) is sufficient for our purposes, no proof of this inequality is publicly available, so we give a complete proof of our concentration inequality here. We restate the lemma for the reader’s convenience.
Lemma (Restatement of Lemma 2.3).
Let $X_{1},\dots,X_{m}$ be a set of independent random variables (often called trials), and let $(\Omega,\Sigma,\Pr)$ be the product space describing all random outcomes of $X_{1},\dots,X_{m}$. Let $\Omega^{*}\subseteq\Omega$ be some set of exceptional outcomes. Let $R_{1},\ldots,R_{n}$ be indicator random variables, each defined as a function of the trials $X_{1},\ldots,X_{m}$, and let $R=R_{1}+\dots+R_{n}$. Suppose that for each non-exceptional outcome $\mathbf{x}=(x_{1},\dots,x_{m})\in\Omega\setminus\Omega^{*}$ of the trials, the following hold for some $\beta,D>0$:
  • $\sum_{i=1}^{n}|W_{i}(\mathbf{x})|\leq D$.
  • $\Pr(\Omega^{*})\leq 1/6$.
Then, for all $\tau\geq 12\sqrt{\pi\beta D}+2\Pr(\Omega^{*})\sup_{\mathbf{x}\in\Omega^{*}}R(\mathbf{x})$,
$\Pr\left(|R-\mathbb{E}[R]|\geq\tau\right)<6\exp\left(-\frac{\tau^{2}}{24\beta D}\right)+\Pr(\Omega^{*}).$
In order to prove Lemma 2.3, we need to introduce some definitions and notation. Given a vector $\mathbf{a}\in\mathbb{R}^{m}$, let $\|\mathbf{a}\|=\sqrt{a_{1}^{2}+\dots+a_{m}^{2}}$. For two outcomes $\mathbf{x}=(x_{1},\dots,x_{m}),\mathbf{y}=(y_{1},\dots,y_{m})\in\Omega$ and a nonzero vector $\mathbf{a}=(a_{1},\dots,a_{m})$, define
$\operatorname{dist}_{\mathbf{a}}(\mathbf{x},\mathbf{y})=\frac{1}{\|\mathbf{a}\|}\sum_{\begin{subarray}{c}i\in\{1,\dots,m\}\\ x_{i}\neq y_{i}\end{subarray}}a_{i}.$
When $\mathbf{a}$ is a unit vector, $\operatorname{dist}_{\mathbf{a}}(\mathbf{x},\mathbf{y})$ gives the $\mathbf{a}$-Hamming distance between $\mathbf{x}$ and $\mathbf{y}$ described by Bruhn and Joos [22]. Then, for an event $\mathcal{A}\subseteq\Omega$ and an outcome $\mathbf{y}\in\Omega$, define
$\operatorname{dist}_{\mathbf{a}}(\mathbf{y},\mathcal{A})=\inf_{\mathbf{x}\in\mathcal{A}}\operatorname{dist}_{\mathbf{a}}(\mathbf{x},\mathbf{y}).$
Next, define
$\operatorname{dist}(\mathbf{x},\mathcal{A})=\sup_{\begin{subarray}{c}\mathbf{a}\in\mathbb{R}^{m}\\ \mathbf{a}\neq\mathbf{0}\end{subarray}}\operatorname{dist}_{\mathbf{a}}(\mathbf{x},\mathcal{A}).$
This notion of distance between an outcome and a set will be useful, as if $\operatorname{dist}(\mathbf{x},\mathcal{A})$ is bounded above, then $\operatorname{dist}_{\mathbf{a}}(\mathbf{x},\mathcal{A})$ is bounded above for every nonzero vector $\mathbf{a}$, in particular for a vector $\mathbf{a}$ whose entries describe relationships between our indicator variables and witnessing trials. Finally, given two events $\mathcal{A}\subseteq\Omega$ and $\mathcal{B}\subseteq\Omega$, define
$\operatorname{dist}(\mathcal{A},\mathcal{B})=\inf_{\mathbf{y}\in\mathcal{B}}\operatorname{dist}(\mathbf{y},\mathcal{A}).$
It is a straightforward exercise to show that $\operatorname{dist}(\mathcal{A},\mathcal{B})=\operatorname{dist}(\mathcal{B},\mathcal{A})$. A general form of Talagrand’s inequality states the following:
Theorem 5.1 ([62]).
Let $\mathcal{A},\mathcal{B}\subseteq\Omega$. If $\operatorname{dist}(\mathcal{A},\mathcal{B})\geq\ell$, then
$\Pr(\mathcal{A})\Pr(\mathcal{B})<\exp\left(-\ell^{2}/4\right).$
With Talagrand’s inequality in hand, we are ready to prove Lemma 2.3.
Proof of Lemma 2.3.
Our proof follows the structure of Mahdian’s proof of the concentration inequality in [49, Section 6]. To account for our exceptional outcomes, we follow a similar strategy to that of Bruhn and Joos in [22, Lemma 9]. We begin with the following claim.
Claim 1.
Let $\mu,\tau>0$ be real numbers, and let
$\mathcal{A}=\{\mathbf{x}\in\Omega\setminus\Omega^{*}\,:\,R(\mathbf{x})\geq\mu\}$ and $\mathcal{B}=\{\mathbf{y}\in\Omega\setminus\Omega^{*}\,:\,R(\mathbf{y})\leq\mu-\tau\}$;
then,
$\Pr(\mathcal{A})\Pr(\mathcal{B})\leq\exp\left(-\frac{\tau^{2}}{4\beta D}\right).$
Proof of claim: We show that $\operatorname{dist}(\mathcal{A},\mathcal{B})\geq\ell$, where $\ell=\tau/\sqrt{\beta D}$. Then, the claim follows from Theorem 5.1. For the sake of contradiction, suppose that some $\mathbf{x}\in\mathcal{A}$ satisfies $\operatorname{dist}(\mathbf{x},\mathcal{B})<\ell$. Then, by definition, for every nonzero $\mathbf{a}\in\mathbb{R}^{m}$, we have $\operatorname{dist}_{\mathbf{a}}(\mathbf{x},\mathcal{B})<\ell$. We construct a nonzero vector $\mathbf{a}\in\mathbb{R}^{m}$ that gives us a contradiction, as follows.
For the outcome $\mathbf{x}$, and for each $R_{i}$ satisfying $R_{i}(\mathbf{x})=1$, we define witness sets $W_{i}(\mathbf{x})$ satisfying the conditions of the lemma, as well as the following auxiliary bipartite graph $H$. Let $H$ have parts $U=\{u_{1},\dots,u_{n}\}$ and $V=\{v_{1},\dots,v_{m}\}$. For each $u_{i}\in U$ and $v_{j}\in V$, add an edge $u_{i}v_{j}$ if and only if $X_{j}\in W_{i}(\mathbf{x})$; that is, if and only if the random variable $X_{j}$ is a witness of the event $R_{i}(\mathbf{x})=1$. Hence, $H$ represents the relationships between the indicator variables $R_{i}$ and their witness sets $W_{i}(\mathbf{x})$. Note also that $H$ is the incidence (Levi) graph of the hypergraph $\mathcal{H}$ whose vertex set is $X_{1},\dots,X_{m}$ and whose edge set is the family of witness sets $W_{i}(\mathbf{x})$. We observe that for each $u_{i}\in U$, $\deg_{H}(u_{i})=|W_{i}(\mathbf{x})|$; that is, $\deg_{H}(u_{i})$ is the number of trials $X_{j}$ that we have chosen to certify that $R_{i}(\mathbf{x})=1$. Similarly, for each $v_{j}\in V$, $\deg_{H}(v_{j})=|\{i:X_{j}\in W_{i}(\mathbf{x})\}|$; in other words, $\deg_{H}(v_{j})$ is the number of indicator variables $R_{i}$ for which $X_{j}$ is a witness. By our lemma’s assumption, $\deg_{H}(v_{j})\leq\beta$ for each $v_{j}\in V$.
We define the vector $\mathbf{a}=(a_{1},\dots,a_{m})$ such that $a_{j}=\deg_{H}(v_{j})$. We observe that
$|E(H)|=\sum_{i=1}^{n}\deg_{H}(u_{i})=\sum_{j=1}^{m}a_{j}=\sum_{i=1}^{n}|W_{i}(\mathbf{x})|\leq D.$
Furthermore, $a_{j}\leq\beta$ for each $j$. Therefore,
$\|\mathbf{a}\|^{2}=\sum_{i=1}^{m}a_{i}^{2}\leq\beta\sum_{i=1}^{m}a_{i}\leq\beta D.$
Now, as $\operatorname{dist}(\mathbf{x},\mathcal{B})<\ell$, we have in particular that $\operatorname{dist}_{\mathbf{a}}(\mathbf{x},\mathcal{B})<\ell$. Thus, there exists $\mathbf{y}\in\mathcal{B}$ such that
$\operatorname{dist}_{\mathbf{a}}(\mathbf{x},\mathbf{y})=\frac{1}{\|\mathbf{a}\|}\sum_{x_{j}\neq y_{j}}a_{j}<\ell.$
As a consequence, $\sum_{x_{j}\neq y_{j}}a_{j}<\ell\|\mathbf{a}\|\leq\ell\sqrt{\beta D}=\tau$.
Now, we claim that there are fewer than $\tau$ indicator variables $R_{i}$ for which $R_{i}(\mathbf{x})=1$ and $R_{i}(\mathbf{y})=0$. Indeed, suppose that $R_{i}(\mathbf{x})=1$ and $R_{i}(\mathbf{y})=0$. This means that the witness set $W_{i}(\mathbf{x})$ contains a variable $X_{j}$ for which $x_{j}\neq y_{j}$. In other words, $u_{i}\in U$ has a neighbor $v_{j}\in V$ for which $x_{j}\neq y_{j}$. The number of vertices $u_{i}\in U$ with a neighbor $v_{j}\in V$ for which $x_{j}\neq y_{j}$ is at most
$\sum_{\begin{subarray}{c}1\leq j\leq m\\ x_{j}\neq y_{j}\end{subarray}}\deg_{H}(v_{j})=\sum_{\begin{subarray}{c}1\leq j\leq m\\ x_{j}\neq y_{j}\end{subarray}}a_{j}<\tau.$
Therefore, the number of indicator variables $R_{i}$ for which $R_{i}(\mathbf{x})=1$ and $R_{i}(\mathbf{y})=0$ is less than $\tau$. As a result, since $R(\mathbf{x})\geq\mu$, it follows that $R(\mathbf{y})>\mu-\tau$, contradicting the assumption that $\mathbf{y}\in\mathcal{B}$. Thus, we conclude that $\operatorname{dist}(\mathcal{A},\mathcal{B})\geq\ell$, and the claim follows. $\blacklozenge$
Now, write $\operatorname{Med}(R)$ for the median of $R=R(\mathbf{x})$ for a randomly sampled $\mathbf{x}\in\Omega$. Then, when $\mathbf{x}\in\Omega$ is randomly sampled,
$\Pr(R\geq\operatorname{Med}(R)\textrm{ and }\mathbf{x}\not\in\Omega^{*})\geq\frac{1}{2}-\frac{1}{6}=\frac{1}{3}.$
Therefore, applying Claim 1 with $\mu=\operatorname{Med}(R)$, we have
$\Pr(R(\mathbf{x})\leq\operatorname{Med}(R)-\tau\text{ and }\mathbf{x}\not\in\Omega^{*})\leq 3\exp\left(-\frac{\tau^{2}}{4\beta D}\right).$
Similarly, if we let $\mu=\operatorname{Med}(R)+\tau$, then
$\Pr(R(\mathbf{x})\geq\operatorname{Med}(R)+\tau\textrm{ and }\mathbf{x}\not\in\Omega^{*})\leq 3\exp\left(-\frac{\tau^{2}}{4\beta D}\right).$
Hence,
$\Pr(|R(\mathbf{x})-\operatorname{Med}(R)|\geq\tau\text{ and }\mathbf{x}\not\in\Omega^{*})\leq 6\exp\left(-\frac{\tau^{2}}{4\beta D}\right).$
As a consequence,
(5.1) $\Pr(|R-\operatorname{Med}(R)|\geq\tau)\leq 6\exp\left(-\frac{\tau^{2}}{4\beta D}\right)+\Pr(\Omega^{*}).$
Now, to show concentration of $R$ around the mean, we show that $|\mathbb{E}[R]-\operatorname{Med}(R)|$ is small. We have
$|\mathbb{E}[R]-\operatorname{Med}(R)|$ $=$ $|\mathbb{E}[R-\operatorname{Med}(R)]|$
$\leq$ $\mathbb{E}|R-\operatorname{Med}(R)|$
$=$ $\sum_{\mathbf{x}\in\Omega\setminus\Omega^{*}}\left(\Pr(\mathbf{x})\cdot|R(\mathbf{x})-\operatorname{Med}(R)|\right)+\sum_{\mathbf{x}\in\Omega^{*}}\left(\Pr(\mathbf{x})\cdot|R(\mathbf{x})-\operatorname{Med}(R)|\right)$
$\leq$ $\sum_{\mathbf{x}\in\Omega\setminus\Omega^{*}}\left(\Pr(\mathbf{x})\cdot|R(\mathbf{x})-\operatorname{Med}(R)|\right)+\Pr(\Omega^{*})\sup_{\mathbf{x}\in\Omega^{*}}R(\mathbf{x})$
We write $\mathbf{1}_{\mathbf{x}\in\Omega\setminus\Omega^{*}}$ for the indicator random variable that equals $1$ when $\mathbf{x}\in\Omega\setminus\Omega^{*}$ and equals 0 when $\mathbf{x}\in\Omega^{*}$. Using the fact that $E[Y]=\int_{0}^{\infty}\Pr(Y\geq y)dy$ for every nonnegative random variable $Y$, we see that
(5.2) $|\mathbb{E}[R]-\operatorname{Med}(R)|$ $\leq$ $\mathbb{E}[|R(\mathbf{x})-\operatorname{Med}(R)|\mathbf{1}_{\mathbf{x}\in\Omega\setminus\Omega^{*}}]+\Pr(\Omega^{*})\sup_{\mathbf{x}\in\Omega^{*}}R(\mathbf{x})$
$=$ $\int_{0}^{\infty}\Pr(|R(\mathbf{x})-\operatorname{Med}(R)|\geq\tau\text{ and }\mathbf{x}\not\in\Omega^{*})d\tau+\Pr(\Omega^{*})\sup_{\mathbf{x}\in\Omega^{*}}R(\mathbf{x})$
$\leq$ $\int_{0}^{\infty}6\exp\left(-\frac{\tau^{2}}{4\beta D}\right)d\tau+\Pr(\Omega^{*})\sup_{\mathbf{x}\in\Omega^{*}}R(\mathbf{x})$
$=$ $6\sqrt{\pi\beta D}+\Pr(\Omega^{*})\sup_{\mathbf{x}\in\Omega^{*}}R(\mathbf{x}).$
Finally, we estimate the probability that $|R-\mathbb{E}[R]|\geq\tau$. Recall our lower bound
$\tau\geq 12\sqrt{\pi\beta D}+2\Pr(\Omega^{*})\sup_{\mathbf{x}\in\Omega^{*}}R(\mathbf{x}).$
With this in hand, we note that if $|R-\mathbb{E}[R]|\geq\tau$, then
(5.3) $|R-\mathbb{E}[R]|\geq\frac{\tau}{2}+6\sqrt{\pi\beta D}+\Pr(\Omega^{*})\sup_{\mathbf{x}\in\Omega^{*}}R(\mathbf{x}).$
Combining (5.2) and (5.3), we have
$|R-\operatorname{Med}(R)|\geq|R-\mathbb{E}[R]|-|\mathbb{E}[R]-\operatorname{Med}(R)|\geq\frac{\tau}{2}.$
Thus, by applying (5.1) with $\tau/2$,
$\Pr(|R-\mathbb{E}[R]|\geq\tau)\leq 6\exp\left(-\frac{\tau^{2}}{24\beta D}\right)+\Pr(\Omega^{*}).$
This completes the proof. ∎
@@ 5. An exceptional variant of a concentration inequality
We aim to apply Lemma 2.3 to concentrate $R$. Therefore, we fix an outcome $\mathbf{x}$ of the Wasteful Coloring Procedure, and we construct a witness set for each random variable $R_{c}$ for which $R_{c}(\mathbf{x})=1$. If $R_{c}(\mathbf{x})=1$, we can construct a witness set $W_{c}(\mathbf{x})$ of size at most one consisting of either $A_{u}$ or $E_{v,c}$. Note that if $W_{c}(\mathbf{x})=\{A_{u}\}$, then $u$ was activated and assigned $c$. Similarly, if $W_{c}(\mathbf{x})=\{E_{v,c^{\prime}}\}$, then $c=c^{\prime}$. Therefore, for each $u\in N_{G}(v)$, $A_{u}$ belongs to at most one witness set $W_{c}(\mathbf{x})$, and for each $c^{\prime}\in L(v)$, $E_{v,c^{\prime}}$ belongs to at most one witness set $W_{c}(\mathbf{x})$. Furthermore, as $\eta d<\ell\leq 8d$ and $\log^{-2}d<\eta$, for sufficiently large $d$ we may conclude,

### v2

@@ 1.2. Vu’s conjecture
In a graph $G$, the codegree of two vertices $u,v$ is the number of distinct neighbors that are common to both $u$ and $v$. Given $\zeta>0$, how large can the chromatic number be in a graph with maximum degree at most $\Delta$ and maximum codegree at most $\zeta\Delta$? Because of its connection to Kahn’s important result on the list chromatic index of linear hypergraphs [39] and, in turn, to the Erdős–Faber–Lovász conjecture [40, 31, 41], Vu [63] famously proposed the following bold conjecture in 2002.
@@
This conjecture remains wide open even if $\chi_{\ell}(G)$ is replaced by the ordinary chromatic number $\chi(G)$. Moreover, even the much weaker statement that $G$ satisfies $\alpha(G)\geq(1/\zeta-\varepsilon)(n/\Delta)$ is still open and was also conjectured by Vu [63]. As mentioned earlier, Conjecture 1.1 is a far-reaching generalization of a theorem of Kahn [39] on the list chromatic index of linear hypergraphs. Vu [63, p. 109] emphasized its significance, noting that: “The bound in [Conjecture 1.1], if true, would be an amazing result. For instance, it would immediately imply a deep theorem of Kahn on the list chromatic index of hypergraphs.” Kahn’s theorem, in turn, implies the asymptotic form of the Erdős–Faber–Lovász conjecture [40, 31, 41]. Recently, Kelly, Kühn, and Osthus [43] confirmed a special case of Conjecture 1.1 that also recovers several of its significant consequences.
@@ 1.3. Proof overview
A key component of our analysis is the interplay between a variant of Talagrand’s inequality (specifically adapted to handle exceptional outcomes) and a Turán-type argument designed to control the vertices of high degree at every step of the nibble process. Concentration tools of this type have been used extensively in recent years [47, 22, 10, 12, 11] so we believe that our approach may be of independent interest. For our proof strategy, we rely on an exceptional variant of Talagrand’s inequality developed by Delcourt and Postle in [27]; see Lemma 2.3 and the discussion afterwards.
Our proof of the concentration of $|X^{\prime}|$ proceeds as follows. Rather than attempting to establish concentration of $|X^{\prime}|$ directly, we first obtain concentration for two auxiliary random variables $|X^{\prime}_{1}|$ and $|X^{\prime}_{2}|$, where $X^{\prime}_{1}$ is the number of vertices $w\in N_{L}(u,c)\cap N_{L}(v,c)$ satisfying a, and $X^{\prime}_{2}$ is the number of vertices $w\in N_{L}(u,c)\cap N_{L}(v,c)$ satisfying a but not b. Since $|X^{\prime}|=|X^{\prime}_{1}|-|X^{\prime}_{2}|$, this reduction allows us to deduce concentration of $|X^{\prime}|$. The following conditions are crucial for bounding the parameter $\beta$ in Lemma 2.3, the variant of Talagrand’s inequality that we apply; $\beta$ captures a Lipschitz-type property of the random variable.
@@
    Let $\Omega^{*}$ be the exceptional event that there exists a color $c^{\prime}$ such that “many” vertices in $X$ are activated and assigned the color $c^{\prime}$. We may assume that this exceptional event does not occur by showing that it has only a small probability of occurring and then applying the exceptional variant of Talagrand’s inequality.
@@ 1.4. Algorithmic considerations
The rest of the paper is organized as follows. In Section 2, we collect some preliminary facts and the main tools we need. In Section 3, we derive Theorem 1.2 and Corollary 1.3 from Theorem 1.4. In Section 4, we describe the Wasteful Coloring Procedure formally and introduce the key “nibble lemma” at the heart of the procedure, which we prove in Section 5. Finally, in Section 6, we iteratively apply the nibble lemma to prove Theorem 1.5.
@@ 2. Preliminaries
In this section, we outline prior known results which are necessary for our proofs. We start by stating the Kővári–Sós–Turán theorem [46].
@@
Suppose that $G$ is a bipartite graph with parts $A$ and $B$ of sizes $m$ and $n$, respectively. Suppose that $G$ has no subset $X\subseteq A$ of size $t$ and no subset $Y\subseteq B$ of size $s$ for which $G[X\cup Y]$ is the complete bipartite graph $K_{s,t}$. Then,
@@
The following is a straightforward corollary of Lemma 2.1 that allows us to apply Kővári–Sós–Turán theorem when the host graph $G$ is not necessarily bipartite. We provide a proof of the corollary for completeness.
@@
Let $G$ be a graph with vertex subsets $A$ and $B$ of sizes $m$ and $n$, respectively, for which $A\cup B=V(G)$ and every edge of $G$ has an endpoint in $A$ and an endpoint in $B$. Suppose that $G$ has no complete bipartite subgraph $K_{s,t}$ with parts $X,Y$, where $X\subseteq A$ is of size $t$ and $Y\subseteq B$ is of size $s$. Then,
@@
Let $H$ be an auxiliary bipartite graph constructed as follows. Let $H$ have disjoint partite sets $A^{\prime}$ and $B^{\prime}$. For each vertex $a\in A$, add a vertex $a^{\prime}$ to $A^{\prime}$, and for each $b\in B$, add a vertex $b^{\prime}$ to $B^{\prime}$. For each pair $(a,b)\in A\times B$, if $ab\in E(G)$, then add an edge $a^{\prime}b^{\prime}$ to $E(H)$. Since each edge of $G$ corresponds to at least one edge of $H$, clearly $e(H)\geq e(G)$. Furthermore, clearly $H$ is $K_{s,t}$-free. Therefore, by Lemma 2.1, we have
$e(G)\leq e(H)<(t-1)^{1/s}nm^{1-1/s}+(s-1)m.$
@@
Next, as mentioned in Section 1.3, we need a concentration inequality that can handle exceptional outcomes.
Lemma 2.3 (Exceptional Talagrand’s inequality [27]).
Let $X_{1},\dots,X_{m}$ be a set of independent random variables (often called trials), and let $(\Omega,\Sigma,\Pr)$ be the product space that describes all random outcomes of $X_{1},\dots,X_{m}$. Let $\Omega^{*}\subseteq\Omega$ be some set of exceptional outcomes. Let $R_{1},\ldots,R_{n}$ be indicator random variables, each defined as a function of the trials $X_{1},\ldots,X_{m}$, and let $R=R_{1}+\dots+R_{n}$. Suppose that for each non-exceptional outcome $\mathbf{x}=(x_{1},\dots,x_{m})\in\Omega\setminus\Omega^{*}$ of the trials, the following hold for some $\beta,r>0$:
@@
  • $|W_{i}(\mathbf{x})|\leq r$ for each $1\leq i\leq n$.
Then, for all $\tau\geq 96\sqrt{r\beta n}+128r\beta+8\Pr(\Omega^{*})n$,
$\Pr\left(|R-\mathbb{E}[R]|\geq\tau\right)<4\exp\left(-\frac{\tau^{2}}{40\beta rn}\right)+4\Pr(\Omega^{*}).$
Lemma 2.3 easily follows from the concentration inequality [47, Theorem 3.4], which is a special case of the linear Talagrand inequality of Delcourt and Postle [27, Theorem 4.4]. The reduction is obtained by defining the verifier function via the witness sets in Lemma 2.3 and observing that, without loss of generality, both $\tau$ and $\mathbb{E}[R]$ are at most $n$. Additionally, letting $\Omega^{*}=\emptyset$ above, we obtain a similar concentration inequality to that of Mahdian [49]. Informally, we apply Lemma 2.3 as follows. Recall the informal description of the Wasteful Coloring Procedure on page 1.3. For each vertex $v$, let $A_{v}$ be the random variable such that $A_{v}=0$ if $v$ is not activated and $A_{v}=\phi(v)\in L(v)$ otherwise. We let $\Omega$ be the product space over the independent random variables $\{A_{v}\}_{v\in V}$. The set $\Omega^{*}\subseteq\Omega$ of exceptional events that we use depends on the random variable that we are interested in concentrating. Typically, we let $\Omega^{*}$ consist of outcomes in which, for a fixed vertex $v$, there is some color $c$ such that many vertices $u\in N(v)$ satisfy $A_{u}=c$. This allows us to control the value of the parameter $\beta$ in Lemma 2.3. The exact application of Lemma 2.3 differs very slightly from the above due to the introduction of equalizing coin flips (see Section 4).
@@
For every $s\geq 2$, there exists a constant $C$ such that the following holds for all $\Delta\geq 2$ and $\zeta\in[\log^{-16s}\Delta,1]$. If $G$ is a graph with maximum degree $\Delta$ and $s$-codegree at most $\zeta\Delta$, then $\chi(G)\leq C\zeta^{1/16s}\Delta$.
@@ 4. The Wasteful Coloring Procedure
We aim to apply Lemma 2.3 to concentrate $R$. Therefore, we fix an outcome $\mathbf{x}$ of the Wasteful Coloring Procedure, and we construct a witness set for each random variable $R_{c}$ for which $R_{c}(\mathbf{x})=1$. If $R_{c}(\mathbf{x})=1$, we can construct a witness set $W_{c}(\mathbf{x})$ of size at most one consisting of either $A_{u}$ for some $u\in N_{G}(v)$ or $E_{v,c}$. Note that if $W_{c}(\mathbf{x})=\{A_{u}\}$, then $u$ was activated and assigned $c$. Similarly, if $W_{c}(\mathbf{x})=\{E_{v,c^{\prime}}\}$, then $c=c^{\prime}$. Therefore, for each $u\in N_{G}(v)$, $A_{u}$ belongs to at most one of the witness sets $W_{c}(\mathbf{x})$, $c\in L(v)$, and for each $c^{\prime}\in L(v)$, $E_{v,c^{\prime}}$ also belongs to at most one of the witness sets $W_{c}(\mathbf{x})$, $c\in L(v)$. Furthermore, as $\eta d<\ell\leq 8d$ and $\log^{-2}d<\eta$, for sufficiently large $d$ we may conclude,